OpenStudy (141823):

Fapronia is driving at a speed of 50 kilometers per hour. How many minutes will it take her to drive 45 miles? Round to the nearest minute.

OpenStudy (141823):

@zepdrix

OpenStudy (141823):

@nincompoop

OpenStudy (anonymous):

Ok so 45 miles is equivalent to 72.4 km

OpenStudy (anonymous):

And her velocity os 50km/hr and as we know since her velocity is constant we can use this formula that t=d/v so 72.4/50 =1.448

OpenStudy (anonymous):

1.448 hrs which is 60min + 0.448*60min=86.88 min which is 87 min

OpenStudy (141823):

thank you so much!

OpenStudy (anonymous):

Anytime :)